# Leveraging Knowledge Graph-Based Human-Like Memory Systems to Solve Partially Observable Markov Decision Processes

### v1

@@ 3.3 The Learning Algorithm
Given the discrete nature of the action spaces for both the baseline and the HumemAI agents, since all of our policies have discrete action spaces, we opt for a value-based RL algorithm (Watkins and Dayan 1992), i.e., DDQN (Hasselt, Guez, and Silver 2016) instead of the vanilla DQN (Mnih et al. 2013), to avoid the overestimation of state-action values. On top of that, we decouple value and advantage in the deep Q-networks (DQN), i.e., dueling network (Wang et al. 2016), to generalize learning across actions. A stochastic gradient descent algorithm, Adam (Kingma and Ba 2015), was used to minimize the temporal difference (TD):
@@
By minimizing this loss, we aim to bring the predicted Q-values closer to the target Q-values, thus improving the policy’s performance over time.
@@
As for $\pi^{\text{mm}}$, out of $N_{\text{observations}}$ observations within a given room, only one of them gets stored in the short-term memory system at a time. This is because each quadruple (observation) requires an action. In the end, all $N_{\text{observations}}$ of them will be moved to the episodic, semantic, or be forgotten, but they are done sequentially, not all at once.
@@
KGE converts a symbolic KG into a list of learnable embeddings. Batch processing for the LSTM requires padding to ensure all samples are of equal length. Memories are ordered to help the LSTM discern the relative differences in qualifier values, such as which memories were stored earlier and which have higher strength values. The output of the LSTM is the final hidden representation, which encapsulates the processed information from the sequence of memories. $max\_val$ is the maximum number of steps in an episode, e.g., 100. $\bm{s}$ helps adjust the influence of different qualifiers based on their significance. $\bm{A}$ performs attention to weigh the contributions of different memory types. The dueling network architecture is integrated within the MLP. It consists of two sub-networks: one for computing the value function and another for computing the advantage values. These sub-networks are then combined to calculate the final Q-values. Figure 3 shows the DQN architecture for the HumemAI agent.
@@ 4 Experiments
We evaluate the baselines with varying sizes of $\bm{h}^{(\tau)}_{t}$. Additionally, for the HumemAI agent, we create two extra variants for ablation: one utilizing only an episodic memory system and the other utilizing only a semantic memory system. This allows us to analyze the individual contributions of each memory type to the overall performance.
We designed the training setup to ensure fair comparison between all agents. The HumemAI agent, trained in two phases, has about 85,000 learnable parameters per phase and undergoes 100 training episodes. The decaying factor semantic memory was set to $0.9$. For parity, baseline agents have 170,000 learnable parameters and 200 training episodes. All other hyperparameters remain constant across agents. We select the best-performing model from training and runs in a separate test environment with a greedy policy. We ran in total of 5 runs to report average performance. Table 1 shows the overall results.
@@ 96 Baseline N/A 155 ($\pm$77)
Table 1: Average test episode rewards for various agent types and capacities. Standard deviations are written in the parentheses (It was run 5 times with random seeds). HumemAI (E) is the HumemAI agent only with an episodic memory system, while (S) is the one only with semantic. The Phase 1 column represents the phase where we learn $\pi^{\text{mm}}$ while $f^{\text{explore}}$ remains fixed with a heuristic. Phase 2 denotes the phase where we learn $\pi^{\text{explore}}$ while $f^{\text{mm}}$ is the neural network learned from Phase 1. One episode consists of 100 steps, where each step involves 10 random questions, resulting in a total of 1,000 questions. Although a theoretical maximum episode reward is 1,000, this is unattainable in practice, as the agent cannot maintain a complete hidden state in its memory throughout the entire episode.
Table 1 shows the superior performance of the HumemAI agent, especially those with both episodic and semantic memory systems. The two-phase training approach proves effective, with Phase 1 optimizing memory management and Phase 2 leveraging this to improve exploration. The HumemAI agents consistently outperform baselines across memory capacities. Interestingly, longer observation histories for baseline agents don’t necessarily improve performance, as seen with the capacity 192 baseline performing worse than the HumemAI agents with capacity 12. This suggests that longer histories may include irrelevant information. Training efficiency is also notable, with the HumemAI agent (capacity 24) requiring 40 minutes in total, while a baseline agent (capacity 48) took about 70 minutes on a single-CPU machine (AMD Ryzen 9 7950X 16-Core Processor). The exact details of the hyperparameters can be found in the supplementary material.
Figure 4 illustrates the attention weights of $\pi^{\text{mm}}$ and $\pi^{\text{explore}}$ over time for the HumemAI agent with capacity 24. In 4(a), we observe a decrease in the agent’s focus on short-term memory as the agent moves between the rooms (observes different sub-graphs), while attention to long-term memory increases. This shift is logical, given that the long-term memory starts empty at the beginning of an episode. Notably, for both $\pi^{\text{mm}}$ and $\pi^{\text{explore}}$, the agent shows a preference for the episodic component of long-term memory. This preference suggests that episodic memories play a crucial role in both memory management and exploration strategies.
@@
Figure 4: Attention weights of $\pi^{\text{mm}}$ and $\pi^{\text{explore}}$, respectively, over time for the HumemAI agent with $capacity=24$
Figure 5 illustrates the agent’s memory systems at the final time step of an episode. It essentially captures information about the hidden belief state. The agent predominantly uses episodic memory for atLocation relations (7 in episodic vs. 4 in semantic), suggesting a preference for timestamped object locations. This strategy likely aids in answering questions by distinguishing between old and new memories. Conversely, map layout information is stored more in semantic memory (8 semantic vs. 5 episodic), reflecting its factual and static nature. Notably, the agent never stores wall-related memories in episodic memory, possibly because the environment never queries wall locations. This memory distribution demonstrates the agent’s ability to adapt its storage strategy based on the nature of the information and its relevance to the task at hand.
Figure 5: An example of the agent’s (the HumemAI agent with $capacity=24$) memory systems, $\bm{M}_{t=99}$. The nodes are colored differently by the memory systems they belong to.
@@ 5 Related Work
Our work stands at the intersection of cognitive science, KG, and RL, and while there has been significant research in each of these areas individually, there is relatively little work that combines them to address the challenges of partially observable environments. Notably, the term “episodic memory” is often used in RL to refer to state-action pairs experienced by the agent, similar to the replay buffer used in DQN (Mnih et al. 2013). However, this usage diverges from our approach, which adheres closely to the cognitive science definition of episodic memory introduced by Tulving et al. (Tulving 1983, 1985; Tulving and Thomson 1973), emphasizing its recency property.
@@
Memory-based RL approaches, such as those by Ming et al. (Meng, Gorbet, and Kulić 2021) and Lin et al. (Lin et al. 2018), add memory components to address partial observability and sample efficiency, respectively. Lin et al. use episodic memory to regularize DQN but differs from our method by not leveraging KGs for interpretability and long-term memory management.
@@ 6 Conclusion
In this paper, we introduced an approach to solving POMDPs by leveraging KG-based human-like memory systems. We developed the Rooms Environment with varying levels of complexity, where agents navigate and answer questions based on partial observations. We compared a baseline agent with a simple memory system to our agent, which incorporates episodic and semantic memory systems. Results show that the HumemAI agent outperforms the baseline, effectively capturing information about the belief state and improving decision-making capabilities.
There is room for improvement in our paper. Potential areas for enhancement include jointly learning memory management and exploration policies, which could yield better performance. Additionally, exploring other function approximators, such as graph neural networks (GNNs), might offer improvements over the currently used LSTM. Transitioning from heuristic-based question-answering functions (the reward function) to learning these functions directly could also be beneficial. Furthermore, applying this method to other POMDP environments could test its generalizability and robustness across different scenarios.

### v2

@@ 3.3 The Learning Algorithm
Given the discrete nature of the action spaces for both the baseline and the HumemAI agents, we opt for a value-based RL algorithm (Watkins and Dayan 1992), i.e., DDQN (Hasselt, Guez, and Silver 2016) instead of the vanilla DQN (Mnih et al. 2013), to avoid the overestimation of state-action values. On top of that, we decouple value and advantage in the deep Q-networks (DQN), i.e., dueling network (Wang et al. 2016), to generalize learning across actions. A stochastic gradient descent algorithm, Adam (Kingma and Ba 2015), was used to minimize the temporal difference (TD):
@@
As for $\pi^{\text{mm}}$, out of $N_{\text{observations}}$ observations within a given room, only one of them gets stored as a short-term memory at a time. This is because each quadruple (short-term memory) requires an action. In the end, all $N_{\text{observations}}$ of them will be moved to the episodic, semantic, or be forgotten, but they are done sequentially, not all at once.
@@
KGE converts a symbolic KG into a list of learnable embeddings. Batch processing for the LSTM requires padding to ensure all samples are of equal length. Memories are ordered to help the LSTM discern the relative differences in qualifier values, such as which memories were stored earlier and which have higher strength values. The output of the LSTM is the final-hidden representation, which encapsulates the processed information from the sequence of memories. $max\_val$ is the maximum number of steps in an episode, e.g., 100. $\bm{s}$ helps adjust the influence of different qualifiers based on their significance. $\bm{A}$ performs attention to weigh the contributions of different memory types. The dueling network architecture is integrated within the MLP. Figure 4 shows the DQN architecture for the HumemAI agent.
@@ 4 Experiments
We evaluate with varying sizes of memory to see how it influences the performance. Additionally, for the HumemAI agent, we create two extra variants for ablation: one utilizing only an episodic memory and the other utilizing only semantic. This allows us to analyze the individual contributions of each memory type to the overall performance.
We designed the training setup to ensure fair comparison between all agents. HumemAI, trained in two phases, has about 85,000 learnable parameters and undergoes 100 training episodes per phase. The decaying factor semantic memory was set to $0.8$. For parity, baseline agents have 170,000 learnable parameters and 200 training episodes. All other hyperparameters remain constant across agents. We select the best-performing model from training and runs in a separate test environment with a greedy policy. We ran in total of 5 runs. Table 1 shows the overall results.
@@ 96 Baseline N/A 155 ($\pm$77)
Table 1: Average test episode rewards (±$\pm\sigma$) for various agent types and capacities, after 5 random seeded runs. Capacity denotes the maximum number of memories stored. HumemAI (E) is only with episodic, while (S) is only with semantic. The best performing agent per capacity is in bold. The Phase 1 column represents the phase where we learn $\pi^{\text{mm}}$ while $f^{\text{explore}}$ remains fixed with a heuristic. Phase 2 denotes the phase where we learn $\pi^{\text{explore}}$ while $f^{\text{mm}}$ is the neural network learned from Phase 1. One episode consists of 100 steps, where each step involves 10 random questions, resulting in a total of 1,000 questions. Although a theoretical maximum episode reward is 1,000, this is unattainable in practice, as the agent cannot maintain a complete hidden state in its memory throughout the entire episode.
Table 1 shows the superior performance of the HumemAI agent, especially those with both episodic and semantic memory. The two-phase training approach proves effective, with Phase 1 optimizing memory management and Phase 2 leveraging this to improve exploration. The HumemAI agents consistently outperform baselines across all memory capacities. Interestingly, longer observation histories for baseline agents do not necessarily improve performance, as seen with the capacity 192 baseline performing worse than the HumemAI agents with capacity 12. This suggests that longer histories may include irrelevant information or confuse the learning process. Training efficiency is also notable, with the HumemAI agent (capacity 48) requiring 40 minutes in total, while a baseline agent (capacity 48) took about 50 minutes on a single-CPU machine (AMD Ryzen 9 7950X 16-Core Processor). The exact details of the hyperparameters can be found in the supplementary material.
Figure 5 illustrates the attention weights of $\pi^{\text{mm}}$ and $\pi^{\text{explore}}$ over time for the HumemAI agent with capacity 48. In 5(a), we observe a decrease in the agent’s focus on short-term memory as the time passes, while attention to long-term memory increases. This shift is logical, given that the long-term memory starts empty at the beginning of an episode. Interestingly, for both $\pi^{\text{mm}}$ and $\pi^{\text{explore}}$, the agent shows a preference change from the episodic to the semantic as the time passes. This indicates that the agent increasingly relies on semantic memory over time, likely because semantic memory encodes generalizable knowledge that becomes more valuable as the episode progresses.
@@
Figure 5: Attention weights of $\pi^{\text{mm}}$ and $\pi^{\text{explore}}$, respectively, over time for the HumemAI agent with $capacity=48$
Figure 6 illustrates the agent’s memory systems at the final time step of a test episode. It essentially captures the summarization of the hidden state, e.g., maze layout, object locations, etc. It can be seen that most of the episodic memories were captured in the beginning of the episode, while most of the semantic memories were captured in the later time horizon, confirming our attention analysis. This memory distribution demonstrates the agent’s ability to adapt its storage strategy based on the nature of the information and its relevance to the task at hand.
Figure 6: An example of the agent’s (HumemAI with $capacity=48$) memory systems, $\bm{M}_{t=99}$. There are 7 short-term, 24 episodic, and 24 semantic memories. The nodes are colored differently by the memory type.
@@ 5 Related Work
Our work stands at the intersection of cognitive science, KG, and RL, and while there has been significant research in each of these areas individually, there is relatively little work that combines them to address the challenges of POMDPs.
One thing to mention is that in RL, the term “episodic memory” is often used to refer to state-action pairs experienced by the agent, similar to the replay buffer used in DQN (Mnih et al. 2013). However, this usage diverges from our approach, which adheres closely to the cognitive science definition of episodic memory introduced by Tulving et al. (Tulving 1983, 1985; Tulving and Thomson 1973), emphasizing its recency property.
@@
Memory-based RL approaches, such as those by Ming et al. (Meng, Gorbet, and Kulić 2021) and Lin et al. (Lin et al. 2018), add memory components to address partial observability and sample efficiency, respectively. Lin et al. use episodic memory to regularize DQN, but differs from our method by not leveraging KGs for interpretability and long-term memory management.
@@ 6 Conclusion
We introduced an approach to solving POMDPs by leveraging KG-based human-like memory systems. We developed the Rooms Environment with varying levels of complexity, where agents navigate and answer questions based on partial observations. We compared a baseline agent with observation history to the HumemAI agent, which incorporates episodic and semantic memory. Results show that our agent that aims to capture the most likely hidden state outperforms the baseline.
There is room for improvement in our paper. For example, jointly learning memory management and exploration policies could yield better performance. Additionally, exploring other function approximators, such as graph neural networks (GNNs), might offer improvements over the currently used LSTM. Transitioning from handcrafted question-answering functions to learning them directly could also be beneficial.